Calculate the pH of each of the following:


1. [OH-] = 3.0 x 10-9


2. 0.024 M HCI


3. 0.022 M NaOH


4. pOH = 3.2

1
Expert's answer
2022-03-30T10:43:04-0400

Solution:

(1):

pOH = −log[OH] = −log(3.0×10−9) = 8.52

pH = 14 − pOH = 14 − 8.52 = 5.48

pH = 5.48


(2):

HCI → H+ + Cl

HCl is a strong acid because it dissociates almost completely.

Therefore,

[H+] = C(HCl) = 0.024 M

pH = −log[H+] = −log(0.024) = 1.62

pH = 1.62


(3):

NaOH → Na+ + OH

NaOH is a strong base because it dissociates almost completely.

Therefore,

[OH] = C(NaOH) = 0.022 M

pOH = −log[OH] = −log(0.022) = 1.66

pH = 14 − pOH = 14 − 1.66 = 12.34

pH = 12.34


(4):

pH = 14 − pOH = 14 − 3.20 = 10.80

pH = 10.80


Answers:

(1): pH = 5.48

(2): pH = 1.62

(3): pH = 12.34

(4): pH = 10.80